The first and second dissociation constants of an acid H2A are 1.0 × 10–5 and 5.0 × 10–10 respectively. The overall dissociation constant of the acid will be ?

Grade:12

2 Answers

Saurabh Koranglekar
2 years ago
Dear student

The overall dissociation constant is equal to the product of first and second dissociation constant. It isKa​=Ka1​×Ka2​=1.0×10−5×5.0×10−10=5.0×10−15

Regards
Vikas TU
2 years ago
H2A ↔ HA− + H+
K1 = [HA−] [H+]/[H2A]
HA−↔ A2− + H+
K1 = [A2−] [H+]/[HA−]
For the reactionH2A ↔ A2− +2 H+
K= [A2−] [H+]^2 / [H2A] = K1 × K2  
 = 1 × 10^−5 × 5  × 10^−10   
= 5 × 10^−15